a chemist has one solution containing 30% insecticide and another solution containing 50% insecticide. How much of each solution should the chemist mix to get 200L of a 42% insecticide?
Answer by mananth(16946)

percent ---------------- quantity
30 ---------------- x
50 ---------------- 200-x
42.00% ---------------- 200
...............
30x+50( 200-x)=42*200
30x+10000-50x=8400
30x-50x =8400-10000
-20x=-1600
/20
x=80 liters 30%
50% ----120 liters